OpenStudy (anonymous):

What is the product of Simplify if possible. sqrt(t)(8 - 3)sqrt(5a)

OpenStudy (anonymous):

\[\sqrt{5at}\]

OpenStudy (anonymous):

OOPS the a wasnt suppose to be in there my bad

OpenStudy (anonymous):

@Anonymous100

OpenStudy (anonymous):

\[\sqrt{5t}\]

OpenStudy (anonymous):

not an option :L

OpenStudy (anonymous):

equation is \[\sqrt{t}(8-3)\sqrt{5}\]

OpenStudy (anonymous):

@Anonymous100

OpenStudy (ajprincess):

\[5\sqrt{5t}\]